OpenStudy (darkigloo):

Find the times t in the open interval 0

OpenStudy (darkigloo):

it also says the particle is at position x=8 at t=0 if that matters

OpenStudy (anonymous):

Do you have any thoughts already?

OpenStudy (darkigloo):

i know that if velocity is positive, the particle is moving left. if velocity is negative, the particle is moving right.

OpenStudy (anonymous):

Are you sure?

OpenStudy (darkigloo):

o.o i thought so

OpenStudy (darkigloo):

oh wait no..

OpenStudy (anonymous):

Didn't you make a typo?

OpenStudy (darkigloo):

if velocity is negative, it moves left. if velocity is positive it moves right.

OpenStudy (anonymous):

It's so good that you finally got it.

OpenStudy (darkigloo):

^.^ but i'm sure there's more to the question.

OpenStudy (anonymous):

OK. When a particle's direction tends to change?

OpenStudy (darkigloo):

i don't know :/

OpenStudy (anonymous):

Assuming you are moving straightly to the right. What if you wanted to change your direction toward the left? What would happen to your speed while you're changing your direction?

OpenStudy (darkigloo):

it would become 0?

OpenStudy (anonymous):

Yes, you now got it. The last step you need to do is to find the zeros of that function.

OpenStudy (darkigloo):

so x=1 and x=2. do oi plug in a number between 0 and 1, 1 and 2, and 2 and 5 to see if it is negative/positive?

OpenStudy (anonymous):

Yes, so we infer immediately that at the two points the direction of the particle will change. Yes, you can plug in numbers to see the direction. But it is not already asked about the direction. Don't worry about it.

OpenStudy (darkigloo):

got it thanks> can you help me with another part of the question?

OpenStudy (anonymous):

I am to get going. Hurry up...

OpenStudy (darkigloo):

ok sorry :/ the average speed of the particle over the interval 0<= t <= 2

OpenStudy (anonymous):

It is too much calculations. You have to obtain the position function x(t) from v(t). [ x(t) = int v(t)dt] Once you get the function x(t) from integration, you'll need to use the average speed function at the interval (0,2). It is defined as \[\Delta x/ \Delta t = (x_2 - x_1)/(t_2 - t_1)\] where t_2 - t_1 = 2 - 0 = 2. You just need to find the respected x values and put them at that function.

OpenStudy (darkigloo):

Alright. Thank you.

OpenStudy (anonymous):

Take care!

OpenStudy (darkigloo):

same to you.